the length of a rectangle is three less than two times the width. the perimeter of the rectangle is 78 inches. find the length and the width of the rectangle
Answer by mananth(16946)

width =x
length = 2x-3
Perimeter =78
Perimeter = 2(x+2x-3)
2(3x-3)=78
6x-6=78
6x=84
/6
x=14 the width
length = 2*14-3=25 inches